# On the maximum ratio between chromatic number and clique number

### v1

@@ 1. Introduction
Ramsey Theory, often described as the investigation of when order arises from chaos, is one of the most influential areas in Graph Theory. The Ramsey number $R(s,t)$ is the minimum $n$ such that every red/blue edge-coloring of the complete graph on $n$ vertices contains either a clique on $s$ vertices colored red or a clique on $t$ vertices colored blue.
The first explicit upper bound for Ramsey numbers was obtained by Erdős and Szekeres [E-Sz] in 1935, who showed that $R(s,t)\leqslant\binom{s+t-2}{s-1}$. Note that in the diagonal case, this shows that $R(k,k)\leqslant 4^{k+o(k)}$. On the other hand, Erdős [Ramsey.lower] established the lower bound $R(k,k)\geqslant 2^{k/2+o(k)}$ in 1947 in one of the first uses of the probabilistic method. Subsequent improvements on the lower bound have only refined it by a constant multiplicative factor (see, e.g., Spencer’s bound obtained via the Lovász Local Lemma [Spencer.LLL]). Since then, one of the major problems in modern combinatorics has been to determine the correct asymptotic behavior of the diagonal Ramsey number $R(k,k)$. More specifically, one could ask whether the limit
@@
exists and, if so, what is its value. This question appears as Problem 1 in [chung], as Problem 77 in https://www.erdosproblems.com/77, and on the webpage https://mathweb.ucsd.edu/~erdosproblems/.
In 1967, Erdős [erdos] asked a seemingly unrelated question. Define $f(n)$ to be the maximum, over all graphs $G$ on $n$ vertices, of the ratio $\frac{\chi(G)}{\omega(G)}$, where $\chi(G)$ denotes the chromatic number of $G$ and $\omega(G)$ the clique number of $G$. Formally, we have
@@
Erdős showed that $f(n)=\Theta(n/(\log n)^{2})$ and asked whether the following limit exists:
@@
This question also appears in [erdos2], as Problem 627 in https://www.erdosproblems.com/627, as Problem 53 in [chung], and on the webpage https://mathweb.ucsd.edu/~erdosproblems/. In this note, we establish a connection between these two problems. In particular, we show that if the following conjecture is true, a solution to the former problem results in a solution to the latter problem.
@@
We remark that Conjecture 1.1 does not seem to appear in any existing work, but it is closely related to the Ramsey Diagonal Conjecture. In Appendix A, we briefly discuss these conjectures. The following theorem establishes a connection between the limits in (1) and (2), should they exist, under the assumption of Conjecture 1.1. Throughout this note, all logarithms are in base 2.
@@
If Conjecture 1.1 holds, and $\lim\limits_{k\to\infty}\frac{\log(R(k,k))}{k}$ exists and is equal to $\mathcal{L}$, then
$f(n)=\big(\mathcal{L}^{2}+o(1)\big)\frac{n}{(\log n)^{2}}.$
In [erdos], Erdős mentioned that by their method, it would be easy to prove that
@@
for constants $c=\frac{1}{4}$ and $C=1$. However, this appears to be a typographical error, as a careful application of their method yields the constants $c=\frac{1}{4}$ and $C=4$. Aided by recent breakthrough developments in Ramsey theory, we can prove the following.
@@
In Section 2, we introduce some notation used throughout the paper and state bounds on $f(n)$, namely Theorems 2.1 and 2.2, establishing a connection between the asymptotic behavior of $f(n)$ and the Ramsey numbers $R(s,t)$. Theorem 1.2 is an immediate consequence of these results, whose proofs appear at the end of Section 2. In Section 3, we establish the numerical bounds of Theorem 1.3. In Appendix A, we discuss versions of the Ramsey Diagonal Conjecture and how they relate to Conjecture 1.1.
@@
In this section, we establish upper and lower bounds sufficient to prove Theorem 1.2. We write $h=o_{n}(t)$ if the functions $t,h\colon{\mathbb{N}}\to\mathbb{R}$ satisfy $\lim\limits_{n\to\infty}\frac{h(n)}{t(n)}=0$. When the variable $n$ is clear from the context, we simply write $h=o(t)$. Let us define
$g(n)\coloneqq\frac{(\log n)^{2}}{n}f(n),\quad L\coloneqq\liminf\limits_{k\to\infty}\;\frac{\log(R(k,k))}{k}\quad\text{ and }\quad M\coloneq\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }}$
The next result establishes a lower bound for the limit inferior $\liminf\limits_{n\to\infty}g(n)$ in terms of $L$, recovering Erdős’ original lower bound $g(n)\geqslant 1/4+o(1)$, and a lower bound for the limit superior $\limsup\limits_{n\to\infty}g(n)$ in terms of $M$.
@@
$\liminf_{n\to\infty}g(n)\geqslant L^{2}\quad\text{ and }\quad\limsup_{n\to\infty}g(n)\geqslant M^{2}.$
The following result establishes upper bounds for the limit superior $\limsup\limits_{n\to\infty}g(n)$.
@@
$\limsup_{n\to\infty}g(n)\leqslant M^{2}.$
Before proceeding with the proofs of these theorems, we observe that
$\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}\leqslant\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }}=M.$ (5)
Fact 2.3.
Under the assumption of Conjecture 1.1, equality holds in (5).
Proof.
For $0<s\leqslant t$, let $k$ be the positive integer such that $(k-1)^{2}<st\leqslant k^{2}$. Let
$D\coloneq\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}.$
Conjecture 1.1 implies that
$R(s,t)\leqslant R(k,k)\leqslant 2^{(D+o_{k}(1))k}$
and, therefore, we have
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{\log(R(k,k))}{k-1}\leqslant D+o_{k}(1)=D+o_{t}(1).\qed$
Combining Theorems 2.1 and 2.2, Fact 2.3 proves Theorem 1.2. We now proceed to prove Theorems 2.1 and 2.2, starting with Theorem 2.1.
@@
$g(n)=\frac{f(n)}{n/(\log n)^{2}}\geqslant\frac{\chi(G)}{\omega(G)}\cdot\frac{(\log n)^{2}}{n}\geqslant\frac{(\log n)^{2}}{\alpha(G)\omega(G)}\geqslant\frac{(\log n)^{2}}{k_{n}^{2}}\geqslant L^{2}+o_{n}(1),$ (7)
where the last inequality is true by (6). Hence, $\liminf\limits_{n\to\infty}g(n)\geqslant L^{2}$.
$\lim\limits_{i\to\infty}\frac{\log(R(s_{i},t_{i}))}{\sqrt{s_{i}t_{i}\ }}=M.$
Then, for $n_{i}\coloneqq R(s_{i}+1,t_{i}+1)-1$, we have $(\log n_{i})^{2}\geqslant s_{i}t_{i}(M^{2}+o_{i}(1))$, since $n_{i}>R(s_{i},t_{i})$. Thus, a similar calculation as in (7) implies that
$g(n_{i})=\frac{f(n_{i})}{n_{i}/(\log n_{i})^{2}}\geqslant M^{2}+o_{i}(1),$
which implies that $\limsup\limits_{n\to\infty}g(n)\geqslant M^{2}$. ∎
Now we focus on proving Theorem 2.2. The proof relies on the following bound on the chromatic number of a graph due to Erdős [erdos], which is obtained by greedily picking maximum independent sets as color classes until few vertices remain, with each one receiving a new color.
@@
If for every subgraph $G^{\prime}$ of $G$ on at least $m_{0}$ vertices we have that $\alpha(G^{\prime})\geqslant r$, then
@@
Finally, we conclude that
$\frac{\chi(G)}{\omega(G)}\leqslant\big(M^{2}+o_{n}(1)\big)\frac{n}{(\log n)^{2}}.\qed$
@@
$R(k,k)\leqslant\left(4e^{-0.14e^{-1}}\right)^{k+o(k)}.$ (9)
This bound follows from the recent breakthrough result by Campos, Griffiths, Morris, and Sahasrabudhe [CGMS], along with its improvement by Gupta, Ndiaye, Norin, and Wei [GNNW]. Precisely, they showed that
@@
If Conjecture 1.1 holds, then Theorem 2.2, Fact 2.3, and (9) imply (4). Indeed,
@@
$\limsup_{n\to\infty}g(n)\leqslant M^{2}\leqslant\max_{x\in(0,1/2]}\varphi(x)^{2}<3.71943.\qed$
Appendix A Ramsey Diagonal Conjecture
One natural question is to understand the growth of the Ramsey Numbers $R(s,t)$ for fixed $s+t$. More precisely, one could ask whether the values of the Ramsey numbers decrease as we move away from the diagonal. A similar question can be asked in the multicolored case. The Diagonal Conjecture poses a positive answer to this question. For two colors, we summarize this conjecture in the following statement.
Conjecture A.1 (Ramsey Diagonal Conjecture (RDC)).
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is widely believed that the Diagonal Conjecture is very difficult to prove. Even in the case $R(t-1,t+1)\leqslant R(t,t)$, there is no relevant progress. Liang, Radziszowski, and Xu [LIANG2019195] discussed several consequences of this conjecture and provided some evidence for its validity.
In [erdos], Erdős used the fact that, given some positive integer $n$, if $\binom{s+t}{t}\geqslant n$, then $st$ is minimized when $st=\lfloor{k/2}\rfloor\lfloor{(k+1)/2}\rfloor$, where $k$ is the smallest integer for which $\binom{k}{\lfloor{k/2}\rfloor}\geqslant n$. The binomial coefficient naturally arises from the first bound on the Ramsey numbers given by Erdős and Szekeres [E-Sz]. This motivates Conjecture 1.1 as a natural bridge to extend Erdős’ proof. We note that Conjecture 1.1 can be stated as a weak version of the following conjecture.
Conjecture A.2 (Multiplicative form of RDC).
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is easy to see that Conjecture A.2 implies Conjecture A.1. This follows from the fact that if $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ and $s_{1}t_{1}\leqslant s_{2}t_{2}$, then $s_{1}+t_{1}\leqslant s_{2}+t_{2}$.

### v2

# A note on the maximum ratio between chromatic number and clique number

@@ 1. Introduction
Ramsey Theory is a fundamental area of Graph Theory concerned with the existence of unavoidable structures in large graphs. The Ramsey number $R(s,t)$ is the minimum $n$ such that every red/blue edge-coloring of the complete graph on $n$ vertices contains either a clique on $s$ vertices colored red or a clique on $t$ vertices colored blue.
While Ramsey [9] proved in 1930 that Ramsey numbers are finite, the first explicit upper bound was obtained by Erdős and Szekeres [3] in 1935, who showed that
$R(s,t)\leqslant\binom{s+t-2}{s-1}$. (1)
Note that in the diagonal case, this shows that $R(k,k)\leqslant 4^{k+o(k)}$. On the other hand, Erdős [4] established the lower bound $R(k,k)\geqslant 2^{k/2+o(k)}$ in 1947 in one of the first uses of the probabilistic method. Subsequent improvements on the lower bound have only refined it by a constant multiplicative factor (see, e.g., Spencer’s bound obtained via the Lovász Local Lemma [10]). Since then, one of the major problems in modern combinatorics has been to determine the correct asymptotic behavior of the diagonal Ramsey number $R(k,k)$. More specifically, one could ask whether the limit
@@
exists and, if so, what is its value [^1] This question appears as Problem 1 in [2], as Problem 77 in https://www.erdosproblems.com/77, and on the webpage https://mathweb.ucsd.edu/~erdosproblems/..
In 1967, Erdős [5] asked a seemingly unrelated question. Define $f(n)$ to be the maximum, over all graphs $G$ on $n$ vertices, of the ratio $\frac{\chi(G)}{\omega(G)}$, where $\chi(G)$ denotes the chromatic number of $G$ and $\omega(G)$ the clique number of $G$. Formally, we have
@@
Erdős showed that $f(n)=\Theta(n/(\log n)^{2})$ and asked [^2] This question also appears in [6], as Problem 627 in https://www.erdosproblems.com/627, as Problem 53 in [2], and on the webpage https://mathweb.ucsd.edu/~erdosproblems/. whether the following limit exists:
@@
In this note, we establish a connection between these two problems. The upper bound on $f(n)$ from [5] combines (1) with the observation that, if $\binom{s+t}{t}\geqslant n$, then
$st\geqslant\lfloor{k/2}\rfloor\lfloor{(k+1)/2}\rfloor,$
where $k$ is the smallest integer satisfying $\binom{k}{\lfloor{k/2}\rfloor}\geqslant n$. This motivates Conjecture 1.1 below as a natural bridge to extend Erdős’ proof.
@@
We were unable to find a previous statement of Conjecture 1.1 in the literature, but it is closely related to the Diagonal Conjecture (Conjecture 1.4; see Section 1.1 for a more detailed discussion). The following theorem establishes a connection between the limits in (2) and (3), should they exist, under the assumption of Conjecture 1.1. Throughout this note, all logarithms are in base 2.
@@
If Conjecture 1.1 holds, and $\lim\limits_{k\to\infty}\frac{\log(R(k,k))}{k}$ exists and is equal to $\ell$, then
$f(n)=\big(\ell^{2}+o(1)\big)\frac{n}{(\log n)^{2}}.$
In [5], Erdős mentioned that by their method, it would be easy to prove that
@@
for constants $c=\frac{1}{4}$ and $C=1$. However, this appears to be a typographical error [^3] Indeed, if $f(n)\leqslant(1+o(1))\frac{n}{(\log n)^{2}}$ is true, then Theorem 2.1 would imply that $R(k,k)\leqslant 2^{k+o(k)}$., as a careful application of their method yields the constants $c=\frac{1}{4}$ and $C=4$.
Aided by recent breakthrough developments in Ramsey theory, we can prove the following.
@@
1.1. Conjecture 1.1 and the Diagonal Conjecture
We note that Conjecture 1.1 is closely related to the following conjecture, which is named the Diagonal Conjecture in [8].
Conjecture 1.4 (Diagonal Conjecture (DC)).
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is widely believed that the Diagonal Conjecture is very difficult to prove. Even in the case $R(t-1,t+1)\leqslant R(t,t)$, there is no relevant progress. Observe that Conjecture 1.1 can be stated as a weak version of the following conjecture.
Conjecture 1.5 (Multiplicative form of DC).
$R(s_{1},t_{1})\leqslant R(s_{2},t_{2}).$
It is easy to see that if $s_{1}\leqslant s_{2}\leqslant t_{2}\leqslant t_{1}$ and $s_{1}t_{1}\leqslant s_{2}t_{2}$, then $s_{1}+t_{1}\leqslant s_{2}+t_{2}$ and therefore Conjecture 1.5 implies Conjecture 1.4.
@@
In Section 2, we introduce some notation used throughout the paper and state bounds on $f(n)$, namely Theorems 2.1 and 2.2, establishing a connection between the asymptotic behavior of $f(n)$ and the Ramsey numbers $R(s,t)$. The proof of Theorem 1.2, which is an immediate consequence of these results, is also given in Section 2. In Section 3, we establish the numerical bounds of Theorem 1.3.
@@
In this section, we establish bounds sufficient to prove Theorem 1.2. Let us define
$g(n)\coloneqq\frac{(\log n)^{2}}{n}f(n),\quad M\coloneq\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }},$
$L\coloneqq\liminf\limits_{k\to\infty}\;\frac{\log(R(k,k))}{k}\quad\text{ and }\quad D\coloneq\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}.$
The following theorems establish more precisely the connection between limits (2) and (3). The first result gives the exactly value of $\limsup g(n)$ in terms of the Ramsey numbers.
@@
$\limsup_{n\to\infty}g(n)=M^{2}.$
The second one establishes a lower bound on $\liminf g(n)$ in terms of the Ramsey numbers.
@@
$\liminf_{n\to\infty}g(n)\geqslant L^{2}.$
Before proceeding to the proof of these results, let us show how to derive Theorem 1.2.
Proof of Theorem 1.2.
The assumption that $\lim\frac{\log R(k,k)}{k}$ exists and equals $\ell$ implies $L=D=\ell$. Thus, by Theorems 2.1 and 2.2, it suffices to prove that $D=M$. To this end, observe that restricting the maximum to $s=t$ in the definition of $M$ yields
$D=\limsup_{k\to\infty}\;\frac{\log(R(k,k))}{k}\leqslant\limsup_{t\to\infty}\;\max_{s\leqslant t}\ \frac{\log(R(s,t))}{\sqrt{st\ }}=M.$ (6)
On the other hand, if $0<s\leqslant t$ and $k$ is the positive integer such that $(k-1)^{2}<st\leqslant k^{2}$, Conjecture 1.1 implies that
$R(s,t)\leqslant R(k,k)\leqslant 2^{(D+o_{k}(1))k}.$
Therefore, we have
$\frac{\log(R(s,t))}{\sqrt{st\ }}\leqslant\frac{\log(R(k,k))}{k-1}\leqslant D+o_{k}(1)=D+o_{t}(1),$
which implies $M\leqslant D$. Combined with (6), we conclude that $D=M$. ∎
We now proceed to prove Theorems 2.1 and 2.2, starting with Theorem 2.2.
@@
$g(n)=\frac{f(n)}{n/(\log n)^{2}}\geqslant\frac{\chi(G)}{\omega(G)}\cdot\frac{(\log n)^{2}}{n}\geqslant\frac{(\log n)^{2}}{\alpha(G)\omega(G)}\geqslant\frac{(\log n)^{2}}{k_{n}^{2}}\geqslant L^{2}+o_{n}(1),$
where the last inequality is true by (7). Hence, $\liminf g(n)\geqslant L^{2}$. ∎
Now we focus on proving Theorem 2.1. The proof relies on the following bound on the chromatic number of a graph from [5], which is obtained by greedily picking maximum independent sets as color classes until few vertices remain, with each remaining vertex receiving a new color.
@@
If $\alpha(G^{\prime})\geqslant r$ for every subgraph $G^{\prime}\subset G$ with at least $m_{0}$ vertices, then
@@
Then, we conclude that
$\frac{\chi(G)}{\omega(G)}\leqslant\big(M^{2}+o_{n}(1)\big)\frac{n}{(\log n)^{2}}.$ (9)
$\lim\limits_{i\to\infty}\frac{\log(R(s_{i},t_{i}))}{\sqrt{s_{i}t_{i}\ }}=M.$
Then, for $n_{i}\coloneqq R(s_{i}+1,t_{i}+1)-1$, we have $(\log n_{i})^{2}\geqslant s_{i}t_{i}\big(M^{2}+o_{i}(1)\big)$, since $n_{i}>R(s_{i},t_{i})$. The definition of $n_{i}$ also implies that there is a graph $G_{i}$ on $n_{i}$ vertices satisfying $\alpha(G_{i})\leqslant s_{i}$ and $\omega(G_{i})\leqslant t_{i}$. Thus we have
$g(n_{i})=\frac{f(n_{i})}{n_{i}/(\log n_{i})^{2}}\geqslant\frac{\chi(G_{i})}{\omega(G_{i})}\cdot\frac{(\log n_{i})^{2}}{n_{i}}\geqslant\frac{(\log n_{i})^{2}}{\alpha(G_{i})\omega(G_{i})}\geqslant\frac{(\log n_{i})^{2}}{s_{i}t_{i}}\geqslant M^{2}+o_{i}(1),$
which together with (9) implies that $\limsup g(n)=M^{2}$.∎
@@
$R(k,k)\leqslant\big(4e^{-0.14e^{-1}}\big)^{k+o(k)}.$ (10)
This bound follows from the recent breakthrough result by Campos, Griffiths, Morris, and Sahasrabudhe [1], along with its improvement by Gupta, Ndiaye, Norin, and Wei [7]. Precisely, they showed that
@@
As seen in the proof of Theorem 1.2, if Conjecture 1.1 holds, then $D=M$. Hence, Theorem 2.1 and (10) imply (5). Indeed,
@@
$\limsup_{n\to\infty}g(n)=M^{2}\leqslant\max_{x\in(0,1/2]}\varphi(x)^{2}<3.71943.\qed$
Acknowledgements
The authors thank Rob Morris for his helpful suggestions that heavily improved the presentation of this paper.